# Mamba-FCS: Joint Spatio-Frequency Feature Fusion, Change-Guided Attention, and SeK Loss for Enhanced Semantic Change Detection in Remote Sensing

### v1

@@ Abstract
Semantic Change Detection (SCD) from remote sensing imagery requires models balancing extensive spatial context, computational efficiency, and sensitivity to class-imbalanced land-cover transitions. While Convolutional Neural Networks excel at local feature extraction but lack global context, Transformers provide global modeling at high computational costs. Recent Mamba architectures based on state-space models offer compelling solutions through linear complexity and efficient long-range modeling. In this study, we introduce Mamba-FCS, a SCD framework built upon Visual State Space Model backbone incorporating, a Joint Spatio-Frequency Fusion block incorporating log-amplitude frequency domain features to enhance edge clarity and suppress illumination artifacts, a Change-Guided Attention (CGA) module that explicitly links the naturally intertwined BCD and SCD tasks, and a Separated Kappa (SeK) loss tailored for class-imbalanced performance optimization. Extensive evaluation on SECOND and Landsat-SCD datasets shows that Mamba-FCS achieves state-of-the-art metrics, 88.62% Overall Accuracy, 65.78% $F_{\mathrm{scd}}$, and 25.50% SeK on SECOND, 96.25% Overall Accuracy, 89.27% $F_{\mathrm{scd}}$, and 60.26% SeK on Landsat-SCD. Ablation analyses confirm distinct contributions of each novel component, with qualitative assessments highlighting significant improvements in SCD. Our results underline the substantial potential of Mamba architectures, enhanced by proposed techniques, setting a new benchmark for effective and scalable semantic change detection in remote sensing applications. The complete source code, configuration files, and pre-trained models will be publicly available upon publication.
Index Terms:
Semantic change detection, Remote sensing imagery, State-space models, Spatial–frequency fusion, Separated Kappa
@@ I Introduction
Furthermore, CD Algorithms can be further divided into supervised and unsupervised approaches. Unsupervised methods such as K-means, ISODATA and graph cut methods aim to identify change labels without the reliance on training with labeled data[9, 10, 11]. However, this tends to limit the performance especially in complex semantic transitions. Supervised methods, on the contrary, leverage labeled data to learn richer representations of these complex transitions in order to gain superior performance[12].
@@
The Attention mechanism, in Transformer Models[20] for incorporating long-range dependencies in sequence modeling, was adopted by Vision Transformers(ViT) for modeling global dependencies in computer vision tasks. These integrate windowed attentions as well as global attention mechanisms for better modeling of global dependencies. These methods and variants have successfully been used in CD Algorithms with competitive results[21]. However, their quadratic complexity renders dense, high‐resolution prediction prohibitively expensive.
@@
Despite these advances, real‑world deployment still suffers from class imbalance, isolation of semantic and change detection branches, and neglect of frequency domain cues. In particular, because certain semantic transitions often occupy a small fraction of pixels in certain datasets, conventional models tend to be biased toward more prevalent classes[27]. Furthermore, although SCD and BCD are naturally interconnected, existing approaches like [28, 29] keep change and semantic decoders isolated, preventing mutual reinforcement and resulting in blurred boundaries and inconsistent bi‐temporal predictions. Moreover, although Fourier‐domain representations are proven to suppress illumination artifacts and emphasize genuine structural differences in BCD[30], existing Mamba frameworks for SCD predominantly focus on spatial features, with limited exploration of frequency-domain fusion benefits.
Furthermore, while the Separated Kappa (SeK) coefficient has become the standard metric for evaluating SCD performance since its introduction on the SECOND benchmark [31], its use as a minimizable loss function remains unexplored.
In this work, we introduce Mamba-FCS, a novel approach that integrates frequency-domain feature fusion and change-guided attention within a Mamba backbone to address limitations of purely spatial processing, and decoder isolation. By incorporating SeK as a loss term, we emphasis semantic consistency within detected change regions. Extensive experiments on the SECOND and LandSat-SCD benchmarks demonstrate substantial improvements in detecting non-salient changes, enhancing recall for rare transitions, and refining semantic boundaries. In summary, our main contributions includes:
@@
  • Proposing the SeK Loss, leveraging the SeK metric as a loss function to directly optimize for semantic consistency within detected change regions, mitigating the impact of class imbalance.
@@ II Related Work
In this section, we review deep learning-based CD methods for high-resolution optical remote sensing imagery, organized into CNN-based, Transformer-based, State Space Model (SSM)-based, and Frequency-Domain-based approaches.
@@ II-A CNN-Based Methods
CNNs have been pivotal in early deep learning approaches for CD due to their robust ability to extract local features. Daudt et al. in [32] proposed the first fully convolutional networks (FCNs) for BCD task, with three variants: FC-Siam-Conc, FC-EF, and FC-Siam-Diff. These models employed a Siamese CNN architecture to process bi-temporal inputs effectively. Subsequent advancements built upon this foundation. For instance, Fang et al. in [33] developed a Siamese-based architecture with a densely connected CNN to enhance feature interaction between bi-temporal images. Similarly, Zhang et al. in [34] introduced multi-level, fine-grained detection by applying deep supervision to differential features extracted at various CNN stages. Shi et al. in [35] further refined this approach by incorporating an attention mechanism to capture more discriminative features. Zhao et al. in [36] investigated the impact of different fusion strategies on BCD performance, proposing a dual encoder-decoder architecture to boost detection accuracy. More recently, Chen et al. in [37] devised an unsupervised sample generation method by swapping image patches, enabling change detection training on single-temporal images.
In contrast to BCD, SCD presents greater challenges but holds significant practical value [38, 39]. Early CNN-based SCD methods adopted a multitask learning framework to simultaneously predict binary change maps and semantic change categories [39]. Ding et al. in [29] proposed an advanced CNN architecture to address the limited information exchange between semantic and change decoders, thereby improving SCD performance.
Despite these advancements, CNN-based models exhibit limitations in CD tasks. CNN architectures suffer from restricted receptive fields, which hinder their ability to capture long-range contextual information essential for detecting semantic changes across large regions. This limitation has driven research toward Transformer-based methods, which offer a global view of the scene [40].
@@ II-B Transformer-Based Methods
With the introduction of Vision Transformers (ViT)[20] in capturing long-range dependencies in images, CD research has increasingly adopted transformer architectures. For BCD, Chen et al. [41] were among the first to apply transformers by converting multi-temporal images into semantic tokens, thereby modeling spatial-temporal relationships to enhance detection accuracy. Bandara et al. in [21] further advanced this by proposing a pure transformer-based Siamese network, which employs a Siamese transformer encoder paired with a multilayer perceptron (MLP) decoder, eliminating the need for convolutional feature extractors.
In the context of SCD, Ding et al.[42] introduced the semantic change transformer, an adaptation of the CSWin Transformer, designed to explicitly learn semantic transitions. This model features a triple encoder-decoder architecture that enhances spatial features and integrates spatio-temporal dynamics with task-specific prior information to reduce learning disparities. Similarly, SMBCNet[43] leverages a transformer-based architecture to perform change detection through semantic segmentation. By integrating a cross-scale enhancement module and a multi-branch change fusion module, SMBCNet effectively captures global information and handles diverse change types. Additionally, STGNet [8] addresses SCD by guiding multitask learning through spatio-temporal semantic interaction, enhancing spatial details and employing a bidirectional guidance module to improve feature extraction in complex scenes.
Despite the advancements brought by transformer-based methods, their computational complexity, primarily due to the quadratic complexity of self-attention mechanisms, poses challenges for large-scale remote sensing datasets. To address this, recent research has explored alternative architectures such as SSMs[22].
Figure 1: Overview of the proposed Mamba-FCS architecture, featuring a VSSM-based encoder and decoder for three tasks- BCD map and SCD maps for $T_{1}$ and $T_{2}$. The encoder processes pre ($I^{T_{1}}$) and post ($I^{T_{2}}$) images through four hierarchical stages (Stage I to Stage IV), extracting multi-scale feature maps. These are fed into the binary change decoder and semantic map decoders, with binary change decoder’s stage-wise outputs guiding the semantic decoders.
@@ II-C SSM-Based Methods
SSMs, particularly the Mamba architecture, have emerged as a promising alternative to Transformers due to their linear computational complexity and ability to model long-range dependencies [23]. In computer vision, the introduction of Visual State Space Model (VMamba) [44] and Vision Mamba [24] has facilitated their adoption for change detection tasks.
For BCD, [28] proposed MambaBCD, which leverages the VMamba architecture to extract comprehensive global spatial features from bi-temporal images. This model uses a change decoder with spatio-temporal relationship modeling, achieving superior performance. Similarly, CDMamba [26] introduces the Scaled Residual ConvMamba (SRCM) block, combining Mamba’s global feature extraction with convolutional layers for enhanced local detail capture. An Adaptive Global Local Guided Fusion (AGLGF) block facilitates bi-temporal feature interaction, improving detection accuracy. MSCNet [45]employs a Mamba-based self-correction network with a spatial-channel interaction fusion architecture, using a momentum update strategy to generate pseudo-labels and correct manual label inaccuracies, enhancing performance on challenging BCD tasks. Additionally, the Iterative Mamba Diffusion Change-Detection Model (IMDCD) [46] integrates Mamba with diffusion models to iteratively refine change detection maps, increasing sensitivity to subtle changes in complex scenes.
For SCD, MambaSCD [28] adapts the Visual Mamba framework to model complex multi-temporal relationships, effectively capturing semantic transitions.
However, above Mamba-based methods primarily operate in the spatial domain, often neglecting frequency-domain information that could suppress illumination artifacts and highlight structural differences in multi-temporal imagery.
@@ II-D Frequency‐Domain–Based Methods
Frequency-domain methods have emerged as a robust approach for remote sensing CD by transforming images from the spatial domain to the frequency domain, typically using the Fourier transform. This transformation highlights structural differences, such as edges and textures, while suppressing illumination artifacts common in multi-temporal imagery [47]. These methods are particularly effective for detecting changes in texture or periodic patterns and for handling multimodal data from diverse sensors, such as optical and synthetic aperture radar (SAR).
For BCD, [48] proposed MFGFNet, which employs multiple global filters in the frequency domain to enhance boundary sharpness and preserve edge information in change regions. By integrating frequency-domain processing with multi-scale feature extraction, MFGFNet improves detection accuracy on datasets like LEVIR-CD. Similarly, the Frequency-Enhanced Mamba for Remote Sensing Change Detection (FEMCD) [30] introduces a difference-guided state-space model (DGSSM) to extract change-related features and a DCT-aided Mamba decoder (DCTMD) to refine minor and texture changes using frequency information, achieving superior BCD performance. SpectMamba [49] further enhances BCD by integrating frequency information with the Mamba architecture to address high-frequency subtle changes and periodic structural changes in multispectral remote sensing images, demonstrating improved performance.
Despite these advancements, to the best of our knowledge, existing literature has not yet extensively explored integrating frequency and spatial information with the Mamba architecture for SCD. Additionally, current Mamba-based methods for SCD often employ isolated decoders, potentially limiting the mutual reinforcement that could benefit both tasks. Moreover, these methods have yet to explore the use of SeK metric as minimizable loss.
Figure 2: Architecture of the Visual State-Space Model (VMamba) backbone, used as the shared encoder. Following a patch-partition layer, four stages of Visual State-Space (VSS) blocks employ 2D selective scanning to progressively downsample spatial resolution (from $H,W$ to $H/32,W/32$) while increasing channel width. Feature shapes extracted from each stage are annotated adjacent to the blocks.
@@ III-A Overview of the Network Architecture
The proposed Mamba-FCS framework utilizes a Siamese network architecture with a shared encoder for processing bi-temporal imagery, followed by three decoders to jointly perform SCD and BCD. As illustrated in Figure 1, the framework processes input image pairs $I^{T_{1}},I^{T_{2}}\in\mathbb{R}^{3\times H\times W}$, where 3 denotes the RGB color channels, and $H$ and $W$ represent the image height and width respectively, corresponding to pre-change and post-change temporal scenes. These bi-temporal inputs are processed by a shared Visual State Space Model (VSSM)-based backbone encoder, denoted as $\mathcal{F}_{encoder}$, which is detailed in Section III-B. This encoder, as seen in figure 2 extracts multi-scale feature maps at four hierarchical levels, providing a robust foundation for subsequent decoding and fusion operations. The feature extraction process is formalized as follows.
@@
These multi-scale feature maps are then fed to the three decoder networks as seen in figure 1. First, our Binary Change Decoder, $\mathcal{F}_{BCD}$, detailed in section III-D, detects changes between the two time steps by fusing corresponding feature maps from $I^{T_{1}}$ and $I^{T_{2}}$ at each stage. The fusion is performed using Joint Spatio-Frequency Feature Fusion Mechanism $\mathcal{F}_{fusion}$ as described in section III-C. As illustrated in figure 4, the Binary Change Decoder employs a hierarchical refinement strategy to generate the final binary change detection output $Y_{BCD}\in\mathbb{R}^{2\times H\times W}$, where the two channels correspond to change and no-change classes. Additionally, the decoder produces intermediate change maps $CM_{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ for $i\in\{1,2,3,4\}$ representing each hierarchical level. This can be formulated as,
@@
The change maps $\{CM_{i}\}_{i=1}^{4}$ will be used in CGA Module in Semantic Change Detection, detailed in the section III-E1.
For SCD, the framework incorporates two dedicated Semantic Map Decoders, $\mathcal{F}_{SCD}^{T_{1}}$ and $\mathcal{F}_{SCD}^{T_{2}}$, one for each time step as detailed in section III-E. These decoders, as illustrated in figure 5, utilize encoder extracted features $\{X_{i}^{T_{j}}\}_{i=1}^{4}$ ($j\in\{1,2\}$) and intermediate change maps $\{CM_{i}\}_{i=1}^{4}$ to enhance semantic segmentation by prioritizing changed regions. The process can be defined as,
@@
Figure 3: The Joint Spatio-Frequency Feature Fusion ($F_{\text{fusion}}$) block, integrated at each decoder stage (left panel), concatenates spatial features $X^{T_{1}}_{i}$, $X^{T_{2}}_{i}$, log-amplitude frequency-domain features $F^{T_{1}}_{i}$, $F^{T_{2}}_{i}$, and the absolute difference map $D_{i}$. These are compressed and refined by a channel-then-spatial CBAM module (right panel) to produce the fused tensor $X^{\text{fused}}_{i}$.
III-B VSSM Backbone Encoder- VMamba
The backbone encoder in the proposed Mamba-FCS framework is based Visual State Space Models (VMamba) [44], a state-of-the-art vision backbone designed for efficient visual representation learning with linear computational complexity. VMamba adapts the Mamba state-space model, originally developed for sequence modeling in natural language processing, to the vision domain, making it particularly suitable for processing bi-temporal imagery due to its ability to capture long-range dependencies with minimal computational overhead [23]. Its promising adaptation to CD tasks is well described in [25].
As depicted in Figure 2, the VMamba encoder employs a hierarchical architecture to process input images. The process begins with a Patch Partition module that divides the input image $I^{T_{j}}$, ($j\in\{1,2\}$) into patches while embedding them into a higher-dimensional feature space. This is followed by four hierarchical stages, each consisting of $L_{i}$ Visual State-Space (VSS) blocks and, except for the first stage, a down-sampling layer. These stages generate feature maps at progressively reduced resolutions ($H/4\times W/4$, $H/8\times W/8$, $H/16\times W/16$, $H/32\times W/32$) with increasing channel depths ($C_{1},C_{2},C_{3},C_{4}$), facilitating the extraction of multi-scale features essential for both BCD and SCD.

### v2

# Mamba-FCS: Joint Spatio-Frequency Feature Fusion, Change-Guided Attention, and SeK Inspired Loss for Enhanced Semantic Change Detection in Remote Sensing

@@ Abstract
Semantic Change Detection (SCD) in remote sensing imagery requires models that integrate extensive spatial context for broad geographic patterns, computational efficiency for large-scale datasets, and sensitivity to class-imbalanced land-cover transitions to detect rare or asymmetric changes. Early SCD approaches relied on Convolutional Neural Networks, which excel in local feature extraction but falter in modeling global spatial context due to limited receptive fields. Transformers mitigate this by capturing long-range dependencies via self-attention, yet their quadratic complexity impairs efficiency on vast remote sensing data. Emerging Mamba architectures, based on state-space models, strike a balance with linear complexity and robust long-range modeling, delivering efficient global context capture and improved performance. In this study, we introduce Mamba-FCS, an SCD framework leveraging a Visual State Space Model backbone, with three key contributions: (1) a Joint Spatio-Frequency Fusion block that integrates log-amplitude frequency-domain features to sharpen edges and mitigate illumination artifacts, (2) a Change-Guided Attention (CGA) module that explicitly bridges the intertwined Binary Change Detection and SCD tasks, and (3) a novel loss function inspired by the Separated Kappa (SeK) metric to optimize for class imbalance. Experiments on the benchmark datasets show that Mamba-FCS consistently outperforms recent state-of-the-art algorithms. Ablation studies indicate that spatio–frequency fusion and CGA mainly sharpen boundaries and suppress hallucinated changes, while the SeK-inspired loss improves minority-class semantics. These results highlight the potential of Mamba-FCS as a scalable and accurate approach for remote sensing change detection. Source code and configuration files are available at https://github.com/Buddhi19/Mamba-FCS.git
@@ I Introduction
Furthermore, CD algorithms can be further divided into supervised and unsupervised approaches. Unsupervised methods such as K-means, ISODATA and graph cut methods aim to identify change labels without the reliance on training with labeled data[32, 1, C_Pérez_2005]. However, this tends to limit the performance especially in complex semantic transitions. Supervised methods, on the contrary, leverage labeled data to learn richer representations of these complex transitions in order to gain superior performance[13].
@@
The Attention mechanism, in Transformer Models[3] for incorporating long-range dependencies in sequence modeling, was adopted by Vision Transformers(ViT) for modeling global dependencies in computer vision tasks. These integrate windowed attentions as well as global attention mechanisms for better modeling of global dependencies. These methods and variants have successfully been used in CD Algorithms with competitive results[2]. However, their quadratic complexity makes dense, high-resolution predictions computationally expensive, limiting scalability for large-scale deployment.
@@
Despite these advances, real‑world deployment still suffers from class imbalance, isolation of semantic and change detection branches, and neglect of frequency domain cues. In particular, because certain semantic transitions often occupy a small fraction of pixels in certain datasets, conventional models tend to be biased toward more prevalent classes[57]. Furthermore, although SCD and BCD are naturally interconnected, existing approaches like [12, 17] keep change and semantic decoders isolated, preventing mutual reinforcement, omitting the inherent inter-dependency from the learning process, and resulting in blurred boundaries and inconsistent bi-temporal predictions. Moreover, although Fourier‐domain representations are proven to suppress illumination artifacts and emphasize genuine structural differences in BCD[48], existing Mamba frameworks for SCD predominantly focus on spatial features, with limited exploration of frequency-domain fusion benefits.
Furthermore, while the Separated Kappa (SeK) coefficient has become the standard metric for evaluating SCD performance since its introduction on the SECOND benchmark [49], its potential as a loss function that support backpropagations for guiding the Deep Learning Model, learnable loss function remains unexplored, representing a clear gap given its ability to leverage the benchmark’s structure for improved learning.
In this work, we introduce Mamba-FCS, a novel approach that integrates frequency-domain feature fusion and change-guided attention within a Mamba backbone to address limitations of purely spatial processing, and decoder isolation. By incorporating SeK inspired loss term, we emphasis semantic consistency within detected change regions. Extensive experiments on the SECOND [49] and LandSat-SCD [51] benchmarks demonstrate substantial improvements in detecting non-salient changes, enhancing recall for rare transitions, and refining semantic boundaries. In summary, our main contributions includes:
@@
  • Proposing the SeK inspired Loss, leveraging the SeK metric as a loss function to directly optimize for semantic consistency within detected change regions, mitigating the impact of class imbalance.
@@ II Related Work
In this section, we review deep learning-based CD methods for high-resolution optical remote sensing imagery, organized into CNN-based, Transformer-based, State Space Model (SSM) based, and frequency-domain-based approaches, and use this overview to highlight the motivation for our study.
@@ II-A CNN-Based Methods
CNNs have been central to early deep learning approaches for CD because they extract local features effectively. Daudt et al. [15] proposed the first fully convolutional networks (FCNs) for BCD, with three variants: FC-Siam-Conc, FC-EF, and FC-Siam-Diff, which use a Siamese CNN architecture to process bi-temporal inputs. Subsequent work built on this foundation. Fang et al. [21] developed a Siamese architecture with a densely connected CNN to strengthen feature interaction between bi-temporal images. Zhang et al. [53] introduced multi-level, fine-grained detection through deep supervision of differential features at several CNN stages. Shi et al. [38] further improved discriminative power by incorporating attention. Zhao et al. [56] studied different fusion strategies for BCD in a dual encoder–decoder architecture that improves detection accuracy. Chen et al. [11] proposed an unsupervised sample generation strategy that swaps image patches so CD models can be trained on single-temporal images.
SCD is more challenging than BCD yet has higher practical value [58, 16]. Early CNN-based SCD methods used multitask learning to predict both binary change maps and semantic change categories [16]. Ding et al. [17] proposed a CNN architecture that increases information exchange between semantic and change decoders and improves SCD performance.
Despite these advances, CNN-based models struggle with long-range context in CD. Their limited receptive fields restrict the capture of large-scale semantic dependencies that are essential for detecting changes across wide regions. This limitation has motivated Transformer-based methods, which provide a global view of the scene [43], and it also motivates architectures that model global context efficiently while enforcing tighter interaction between binary and semantic change branches.
@@ II-B Transformer-Based Methods
The introduction of Vision Transformers (ViT) [3] for long-range dependency modeling in images has led to growing use of transformer architectures in CD. For BCD, Chen et al. [7] were among the first to apply transformers by converting multi-temporal images into semantic tokens and modeling spatial–temporal relationships to improve detection accuracy. Bandara et al. [2] then proposed a pure transformer-based Siamese network with a Siamese transformer encoder and an MLP decoder, removing the need for convolutional feature extractors.
For SCD, Ding et al. [18] introduced the semantic change transformer, an adaptation of the CSWin Transformer that explicitly learns semantic transitions. It uses a triple encoder–decoder architecture that enhances spatial features and integrates spatio-temporal dynamics with task-specific priors to reduce learning disparities. SMBCNet [22] uses a transformer-based architecture for change detection through semantic segmentation, combining a cross-scale enhancement module with a multi-branch change fusion module to capture global information and handle diverse change types. STGNet [44] further improves SCD by guiding multitask learning through spatio-temporal semantic interaction, refining spatial details and using a bidirectional guidance module to enhance feature extraction in complex scenes.
Transformer-based methods significantly advance CD, yet their quadratic self-attention complexity is problematic for large-scale remote sensing datasets. This has driven interest in alternative architectures such as SSMs [25].
Figure 1: Overview of the proposed Mamba-FCS architecture. A Siamese encoder processes the pre-change image $I^{T_{1}}$ and post-change image $I^{T_{2}}$ through four stages (Stage I–IV) to extract multi-scale features. These features feed a central binary change decoder, which predicts the binary change map $Y_{\mathrm{BCD}}$, and two symmetric semantic decoders that output the semantic maps $Y^{T_{1}}$ and $Y^{T_{2}}$. Arrows of type (a) and (b) denote multi-scale skip connections from the pre-change and post-change encoder branches, respectively, to their corresponding decoder stages, (c) indicates within-branch propagation in the semantic decoders, and (d) marks change-guidance connections, where features from the binary decoder are injected into both semantic decoders.
@@ II-C SSM-Based Methods
SSMs, and especially the Mamba architecture, are promising alternatives to Transformers because they offer linear computational complexity while still modeling long-range dependencies [24]. In computer vision, Visual State Space Model (VMamba) [29] and Vision Mamba [60] have enabled SSMs to be used for change detection.
For BCD, MambaBCD [12] uses VMamba to extract global spatial features from bi-temporal images and employs a change decoder with spatio-temporal relationship modeling that yields strong performance. CDMamba [55] introduces the Scaled Residual ConvMamba (SRCM) block, which combines Mamba’s global feature extraction with convolutional layers to strengthen local detail capture, and an Adaptive Global Local Guided Fusion (AGLGF) block that improves bi-temporal feature interaction. MSCNet [41] uses a Mamba-based self-correction network with a spatial–channel interaction fusion architecture and a momentum update strategy to generate pseudo-labels and correct noisy annotations, improving performance on challenging BCD tasks. The Iterative Mamba Diffusion Change-Detection Model (IMDCD) [28] couples Mamba with diffusion models to refine change maps iteratively and increase sensitivity to subtle changes in complex scenes.
For SCD, MambaSCD [12] adapts VMamba to model complex multi-temporal relationships and effectively capture semantic transitions.
Most existing Mamba-based methods operate only in the spatial domain and rarely exploit frequency-domain information, which could suppress illumination artifacts and emphasize structural differences in multi-temporal imagery.
@@ II-D Frequency‐Domain–Based Methods
Frequency-domain methods transform images from the spatial domain using the Fourier transform and have become a robust option for remote sensing CD. The frequency representation highlights structural differences such as edges and textures and suppresses illumination artifacts that are common in multi-temporal imagery [31]. These methods are effective for detecting texture changes or periodic patterns and for handling multimodal data from sensors such as optical imagers and Synthetic Aperture Radar (SAR) imagery.
For BCD, MFGFNet [52] uses multiple global filters in the frequency domain to enhance boundary sharpness and preserve edges in change regions. By combining frequency-domain processing with multi-scale feature extraction, MFGFNet improves accuracy on datasets such as LEVIR-CD [5]. FEMCD [48] introduces a difference-guided state-space model (DGSSM) to extract change-related features and a DCT-aided Mamba decoder (DCTMD) that refines minor and texture changes using frequency cues, achieving strong BCD performance. SpectMamba [19] further integrates frequency information with the Mamba architecture to handle high-frequency subtle changes and periodic structural changes in multispectral remote sensing images and obtains improved results.
To the best of our knowledge, there is still limited work on integrating frequency and spatial information within the Mamba architecture for SCD. Existing Mamba-based SCD methods often use separate decoders, which can weaken mutual reinforcement between binary and semantic branches. They also do not use the SeK metric as a minimizable loss.
These observations motivate our Mamba-FCS framework. It performs joint spatial–frequency fusion within a Mamba encoder–decoder, introduces change-guided semantic decoders that share information with the binary branch, and employs a SeK-inspired loss to better capture rare semantic transitions. Unlike existing Transformer and Mamba-based SCD models that mainly use spatial features, keep binary and semantic decoders loosely coupled, and rely on generic losses for class imbalance, Mamba-FCS is designed to exploit frequency-domain cues, enforce semantic consistency through change-guided interaction from the binary change map, and directly optimize a SeK-inspired objective under severe class imbalance.
Figure 2: Architecture of the Visual State-Space Model (VMamba) backbone, serving as the shared encoder. Following an initial patch-partition layer, the encoder comprises four stages $i=1,2,3,4$, each with $L_{i}$ Visual State-Space (VSS) blocks. These blocks utilize 2D selective scanning to progressively downsample the spatial resolution (from $H\times W$ to $H/32\times W/32$) while expanding the channel depth. Feature dimensions extracted from each stage are annotated adjacent to the corresponding blocks.
@@ III-A Overview of the Network Architecture
The proposed Mamba-FCS framework utilizes a Siamese network architecture with a shared encoder for processing bi-temporal imagery, followed by three decoders to jointly perform SCD and BCD. As illustrated in Figure 1, the framework processes input image pairs $I^{T_{1}},I^{T_{2}}\in\mathbb{R}^{3\times H\times W}$, where 3 denotes the RGB color channels, and $H$ and $W$ represent the image height and width respectively, corresponding to pre-change and post-change temporal scenes. These bi-temporal inputs are processed by a shared Visual State Space Model (VMamba) based backbone encoder, denoted as $\mathcal{F}_{encoder}$, which is detailed in Section III-B. This encoder, as seen in figure 2 extracts multi-scale feature maps at four hierarchical levels, providing a robust foundation for subsequent decoding and fusion operations. The feature extraction process is formalized as follows.
@@
These multi-scale feature maps are then fed to the three decoder networks as seen in figure 1. First, our Binary Change Decoder, $\mathcal{F}_{BCD}$, detailed in section III-D, detects changes between the two time steps by fusing corresponding feature maps from $I^{T_{1}}$ and $I^{T_{2}}$ at each stage. The fusion is performed using our novel Joint Spatio-Frequency Feature Fusion Mechanism $\mathcal{F}_{fusion}$ as described in section III-C. As illustrated in figure 4, the Binary Change Decoder employs a hierarchical refinement strategy to generate the final binary change detection output $Y_{BCD}\in\mathbb{R}^{2\times H\times W}$, where the two channels correspond to change and no-change classes. Additionally, the decoder produces intermediate change maps $CM_{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ for $i\in\{1,2,3,4\}$ representing each hierarchical level. This can be formulated as,
@@
The change maps $\{CM_{i}\}_{i=1}^{4}$ will be used in our CGA Module in Semantic Change Detection, detailed in the section III-E1.
For SCD, the framework incorporates two dedicated Semantic Map Decoders, $\mathcal{F}_{SCD}^{T_{1}}$ and $\mathcal{F}_{SCD}^{T_{2}}$, one for each time step as detailed in section III-E. These decoders, as illustrated in figure 5, utilize encoder extracted features $\{X_{i}^{T_{j}}\}_{i=1}^{4}$ ($j\in\{1,2\}$) and intermediate change maps $\{CM_{i}\}_{i=1}^{4}$ extracted from Binary Change Detector to enhance semantic segmentation by prioritizing changed regions (The information flow can be seen in figure 1). The process can be defined as,
@@
Figure 3: (a) The Joint Spatio-Frequency Feature Fusion ($F_{\text{fusion}}$) block, embedded within each decoder stage $i$, concatenates spatial features $X^{T_{1}}_{i}$ and $X^{T_{2}}_{i}$, log-amplitude frequency-domain features $F^{T_{1}}_{i}$ and $F^{T_{2}}_{i}$, and the absolute difference map $D_{i}$ to form $X_{i}^{cat}$. (b) The CBAM Refinement Module compresses and refines $X_{i}^{cat}$ through a Convolutional Block Attention Module (CBAM), yielding the fused output tensor $X^{\text{fused}}_{i}$.
III-B Visual State Space Model Based Backbone Encoder- VMamba
The backbone encoder in the proposed Mamba-FCS framework is based on Visual State Space Models (VMamba) [29], a state-of-the-art (SOTA) vision backbone designed for efficient visual representation learning with linear computational complexity. VMamba adapts the Mamba state-space model, originally developed for sequence modeling in natural language processing, to the vision domain, making it particularly suitable for processing bi-temporal imagery due to its ability to capture long-range dependencies with minimal computational overhead [24]. Its promising adaptation to CD tasks is well described in [12].
As depicted in Figure 2, the VMamba encoder employs a hierarchical architecture to process input images. The process begins with a Patch Partition module that divides the input image $I^{T_{j}}$, ($j\in\{1,2\}$) into patches while embedding them into a higher-dimensional feature space. This is followed by four hierarchical stages, each consisting of $L_{i}$ Visual State-Space (VSS) blocks and, except for the first stage, a down-sampling layer. These stages generate feature maps at progressively reduced resolutions ($H/4\times W/4$, $H/8\times W/8$, $H/16\times W/16$, $H/32\times W/32$) with increasing channel depths ($C_{1}<C_{2}<C_{3}<C_{4}$), facilitating the extraction of multi-scale features essential for both BCD and SCD.